# On Thermalization in A Nonlinear Variant of the Discrete NLS Equation

### v1

@@ II Toy Model System & Thermodynamic Formalism
Figure 1: Phase diagram for the toy model with (a) $D=0.25$ and (b) $D=2$. In (a), the temperature curves for $\beta=0$ and $\beta=\infty$ are given by $h=a^{2}/2$ (blue line, TIO blue circles) and $h=a^{2}/8$ (black line, TIO magenta circles), respectively. In both panels, red circles indicate initial states characterized by a decay exponent $\gamma=2$ in the probability distribution of excursion times, marking the ergodic–nonergodic crossover. The orange dashed line shows the maximum energy achievable with a constant norm distribution, $h_{\text{max}}(A_{\ell}=a)=\tfrac{a^{2}}{4}(1+2D)$. The black solid line, $h_{\text{min}}(A_{\ell}=a)=\tfrac{a^{2}}{4}(1-2D)$, is the corresponding minimum for the initial states with a homogeneous norm and coincides with the ground state $h(\beta=\infty)$ for $D=0.25$ in (a). In (b), this $h_{\text{min}}(A_{\ell}=a)$ line is not the true ground state, as heterogeneous norm distributions allow energies that become unbounded from below in the thermodynamic limit for $D>0.5$. While the TIO is only applicable in the cyan region of panel (a), each of the different regimes and their region of applicability are discussed in the text.
@@
These compactly supported states are constructed as follows. After we prepare the high-energy initial state as in Eq.(14), we set $A_{0}=\text{max}(A_{\ell})/m$, and then additionally excite $m$ adjacent sites with the amplitude $A_{0}/m$ in the middle of the lattice with $b_{S}=\sqrt{A_{0}}e^{i\phi_{S}}$ where $S=\{\ell_{0},\ldots,\ell_{0}+m-1\}$. We set $\phi_{S}=S\Delta\phi=S\pi/2$ unless it is a test run for the low-energy in-phase structures, where $\Delta\phi$ is set to zero. For $m=2$, we choose $\ell_{0}=N/2$, and for $m=3$, $\ell_{0}=N/2-1$. The initial state for the compacton tests can be described as follows
@@
Fig. 1 shows the phase diagrams of energy density $h$ versus norm density $a$ for $D=0.25$ and $D=2$. According to equilibrium statistical mechanics [44, 2] the Hamiltonian has to be bounded from below to have a convergent partition function. In both panels, the solid black line represents the minimum energy reached by a constant norm density, $h_{\mathrm{min}}(a_{\ell}=a)=\frac{a^{2}}{4}(1-2D)$ (see Eq. (A.2) for details). For $D=0.25<0.5$, this curve is convex and corresponds to the true ground-state line, meaning it is size-independent and the region below is inaccessible to any initial state. For $D=2$, the same equation becomes concave and no longer represents the ground state. In this case, by heterogeneously redistributing the norm, the energy can be reduced up to $h_{\min}(a)=-7a^{2}N/44$ (see Eq. (26) below). The latter is the minimum for $D=2$ [35] that is reached by only exciting three sites with $A_{\ell}=A_{\ell+2}=3,A_{\ell+1}=5$ (for $D=2$, the minimum energy is achieved with exciting only nearest three sites with a 3-5-3 proportionality of the respective square moduli) and normalizing the lattice to have the desired $a$, as described in the previous paragraph. In the thermodynamic limit $N\to\infty$, the minimum energy density becomes unbounded from below, and no global ground-state line exists. Hence, note that the gray area in Fig 1(b) is accessible and non-Gibbs thermalizable in its nature, while the gray area in Fig.1(a) is inaccessible.
@@
Figure 2: Finite time averages $q(T)$ for $a=0.25$ with (a) $D=0.25$ for $h=0.02,0.03125,0.05,0.07,0.1$, and (b) $D=2$ for $h=0.02,0.03125,0.1,1,2$.
@@ III.1 Finite Time Averages
Figure 3: The probability density function of the initial states for $a=0.25$ with (a) $D=0.25$, (b) $D=2$ for different energy densities $h$. The black dashed lines in both panels refer to a power law decay with an exponent $\gamma=2$ ($\sim 1/{\tau_{+}}^{2}$), to point out the results that has $\gamma\leq 2$ (nonergodic), and results that has $\gamma>2$ (ergodic).
@@ III.2 PDF of Excursion Times
As in the previous section, we choose the local norm densities, $a_{\ell}=|b_{\ell}|^{2}$, as our local observables. The diagnostic developed herein proceeds by calculating the mean norm density $a$ for the entire system and then tracking each $a_{\ell}(t)$. We count the instances when $a_{\ell}(t)$ crosses the equilibrium value (or so-called the ensemble average of the observables $\langle a_{\ell}\rangle_{\Gamma}=a$) and measure the time intervals between these piercings, which are called excursion times [9, 31]. Specifically, we classify the excursion times $\tau$ based on whether $a_{\ell}>a$ or $a_{\ell}<a$, denoting them as $\tau_{+}$ and $\tau_{-}$, respectively. We then determine the probability distribution function (PDF) for $\tau_{+}$ and $\tau_{-}$. To compute the $PDF(\tau_{\pm})$, we bin the values of $\tau_{\pm}$ and normalize by the total number of events.
@@
where $PDF(\tau)$ denotes either $PDF(\tau_{+})$ or $PDF(\tau_{-})$, $\tau_{0}>0$ is a lower cutoff (a minimal excursion time), and the power-law form is assumed for the tail of $PDF(\tau)$. If the power-law decay exponent $\gamma\leq 2$, the mean of $\tau$ diverges, which signifies infinitely large excursion times. Thus, we expect $\gamma\leq 2$ for a nonergodic system (see [31] for details).
@@ III.3 Mechanism for few-site localization
$H(x,y)=\frac{1}{2}x^{2}+\frac{1}{4}y^{2}-D\,xy\cos{(2\Delta\phi)}$
For the staggered compacton ($\Delta\phi=\pi/2$), extremize
which gives the stationarity system
$\partial_{x}\mathcal{L}$ $:$ $x+Dy+2\lambda=0,$ (22)
$\partial_{y}\mathcal{L}$ $:$ $\tfrac{1}{2}y+Dx+\lambda=0,$ (23)
$2x+y=A.$ (24)
Solving yields
and the corresponding energy
$H_{3,\mathrm{stag}}(D)=\frac{1-2D^{2}}{4(3-4D)}\,A^{2}.$ (25)
An analogous three–site calculation for the in–phase configuration ($\Delta\phi=0$) yields
$x=\frac{1+D}{4D+3}\,A,\qquad y=\frac{1+2D}{4D+3}\,A,$
and
$H_{3,\mathrm{in}}(D)=\frac{1-2D^{2}}{4(4D+3)}\,A^{2},$ (26)
so that for $D=2$ one finds $H_{3,\mathrm{in}}=-7A^{2}/44$, which is the global minimum at fixed $A$; cf. also the associated discussions of [29] and [35].
From Eqs.(20),(21),(25),(26); we obtain that
@@
Specifically, for $D=0.25$:
$H_{1}=\tfrac{1}{4}\,A^{2}=0.25\,A^{2},$
$H_{2,\mathrm{in}}=\tfrac{3}{32}\,A^{2}\approx 0.094\,A^{2},\hskip 10.00002ptH_{2,\mathrm{stag}}=\tfrac{5}{32}\,A^{2}\approx 0.16\,A^{2},$
$H_{3,\mathrm{in}}=\tfrac{7}{128}\,A^{2}\approx 0.055\,A^{2},\hskip 10.00002ptH_{3,\mathrm{stag}}=\tfrac{7}{64}\,A^{2}\approx 0.1\,A^{2}.$
Hence $H_{1}$ has the highest energy as stated in Eq.(27). For large $H$ at fixed $A$, the dynamics select the constrained maximizer, i.e., the 1-site state for $D<1$ [35], consistent with our $a_{\ell}(t)$ results (see Figs. 4, 6).
Setting $D=2$ in Eqs. (20), (21), and (25), we find
$H_{1}=\tfrac{1}{4}\,A^{2},\quad H_{2,\mathrm{in}}=-\tfrac{1}{8}\,A^{2},\quad H_{2,\mathrm{stag}}=\tfrac{3}{8}\,A^{2},\quad$
$H_{3,\mathrm{in}}=-\tfrac{7}{44}\,A^{2},\quad H_{3,\mathrm{stag}}=\tfrac{7}{20}\,A^{2},$
where $H_{2,\mathrm{stag}}=0.375A^{2}$ has the highest energy, closely followed by $H_{3,\mathrm{stag}}=0.35A^{2}$, in the order as in Eq.(28). Hence, for $D>1$, the constrained maximizer in the staggered family is the two-site staggered compacton [35], which dominates at large $H$ for fixed $A$. It is thus relevant to appreciate that the predominant dynamical observation in the case of $D=2$ involves such a 2-site staggered configuration, rather than the three-site states which are not generic in this case. Indeed, in our tests the latter appear only when the initial data excite two or more adjacent sites with nearly equal phases (see Eq.(15) as an example initial state used for such observation).
@@
Figure 5: High-energy localization on two and three sites for $D=2$. Local norm densities $a_{\ell}$ are color-coded in time for $a=0.25,h=1$. We present the whole lattice in top figure. We zoom to a 3-site, in-phase localization in the middle panel. We zoom to the dominant 2 site staggered localization in bottom. The initial state is prepared according to Eq.(15) setting $a=0.25,x=19,c=21$.
Figure 6: Nonergodic dynamics exhibits a one-site compacton for $D<1$ and a two-site staggered compacton for $D>1$. Local norm densities are shown as a function of time for $a=0.25$: (a) $D=0.75$, $h=0.57$ and (b) $D=1.25$, $h=1.25$. The initial states are prepared using Eq. (14) with the same random realization of $\eta_{\ell}$ and $\epsilon_{\ell}$, using $(c,x)=(4,12)$ in (a) and $(c,x)=(10,18.13)$ in (b). For clarity, we plot only a subset of lattice sites (a zoom in $\ell$) to highlight the localization pattern.
@@ IV Discussion
For $D=0.25$, the system shows predominantly 1-site localization (see Fig. 4), whereas for $D=2$, we observed 2-site and 3-site stable localization (see Fig. 5). The 2-site (staggered) structures are generically favored, although, as is shown in the figure, an initialization of a 3-site in-phase excitation yields a structure that is quite long-lived, suggesting connections with the work of [29] on energy minimizers. These findings are consistent with previous work [35], which identified multi-site compactons in similar nonlinear systems. The behavior of the toy model for different values of $D$ correlates well with the theoretical predictions of localization patterns based on energy maximizers.
@@ Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 414984028 – SFB 1404 FONDA and supported through the DFG Walter Benjamin Programme (Grant No. 506198590) (Y.K.). We acknowledge support from the Ministry of Science, Technological Development and Innovation of the Republic of Serbia, Grant No. 451-03-136/2025-03/200017 and 451-03-137/2025-03/ 200124. This material is based upon work supported by the US National Science Foundation under Grant No. PHY-2110030, PHY-2408988 and DMS-2204702 (P.G.K.). This research was partly conducted while P.G.K. was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP) and the Sydney Mathematical Research Institute (SMRI) of the University of Sydney. This work was also supported by the Simons Foundation [SFI-MPS-SFM-00011048, P.G.K.].
@@
The data supporting the findings of this study are available from the corresponding author upon reasonable request.
@@ Appendix A Phase Diagram for Toy Model
A.1 Infinite temperature, $\beta\rightarrow 0$
For the norm density, we have:
$a=-\frac{1}{\beta N}\frac{\partial\ln{Z}}{\partial\alpha}=-\frac{1}{\beta\lambda_{0}}\frac{\partial\lambda_{0}}{\partial\alpha}$ (29)

### v2

@@ II Toy Model System & Thermodynamic Formalism
Figure 1: Phase diagram for the toy model: (a) $D=0.25$ and (b) $D=2$. In (a), the $\beta=0$ and $\beta\rightarrow\infty$ curves correspond to $h=a^{2}/2$ (blue line, TIO blue circles) and $h=a^{2}/8$ (black line, TIO magenta circles). Red circles denote initial states with excursion-time exponent $\gamma=2$. The orange dashed line is the homogeneous-norm upper bound $h_{\max}(A_{\ell}=a)=\tfrac{a^{2}}{4}(1+2D)$, and the black solid line the corresponding lower bound $h_{\min}(A_{\ell}=a)=\tfrac{a^{2}}{4}(1-2D)$; for $D=0.25$ this coincides with $h(\beta\rightarrow\infty)$. For $D>0.5$ [panel (b)], heterogeneous norm distributions allow energies unbounded from below in the thermodynamic limit, so $h_{\min}(A_{\ell}=a)$ is not the ground state. The TIO approach is applicable only within the cyan region of panel (a).
@@
These compactly supported states are constructed as follows. After we prepare the high-energy initial state as in Eq.(14), we set $A_{0}=\text{max}(A_{\ell})/m$, and then additionally excite $m$ adjacent sites in the middle of the lattice with $b_{S}=\sqrt{A_{0}}e^{i\phi_{S}}$ where $S=\{\ell_{0},\ldots,\ell_{0}+m-1\}$. We set $\phi_{S}=S\Delta\phi=S\pi/2$ unless it is a test run for the low-energy in-phase structures, where $\Delta\phi$ is set to zero. For $m=2$, we choose $\ell_{0}=N/2$, and for $m=3$, $\ell_{0}=N/2-1$. The initial state for the compacton tests can be described as follows
@@
Fig. 1 shows the phase diagrams of energy density $h$ versus norm density $a$ for $D=0.25$ and $D=2$. According to equilibrium statistical mechanics [44, 2] the Hamiltonian has to be bounded from below to have a convergent partition function. In both panels, the solid black line represents the minimum energy reached by a constant norm density, $h_{\mathrm{min}}(a_{\ell}=a)=\frac{a^{2}}{4}(1-2D)$ (see Eq. (40) for details). For $D=0.25<0.5$, this curve is convex and corresponds to the true ground-state line, meaning it is size-independent and the region below is inaccessible to any initial state. For $D=2$, the same equation becomes concave and no longer represents the ground state. In this case, by heterogeneously redistributing the norm, the energy can be reduced up to $h_{\min}(a)=-7a^{2}N/44$ (see Eq. (25) below). The latter is the minimum for $D=2$ [35] that is reached by only exciting three sites with $A_{\ell}=A_{\ell+2}=3,A_{\ell+1}=5$ (for $D=2$, the minimum energy is achieved with exciting only nearest three sites with a 3-5-3 proportionality of the respective square moduli) and normalizing the lattice to have the desired $a$, as described in the previous paragraph. In the thermodynamic limit $N\to\infty$, the minimum energy density becomes unbounded from below, and no global ground-state line exists. Hence, note that the gray area in Fig 1(b) is accessible and non-Gibbs thermalizable in its nature, while the gray area in Fig.1(a) is inaccessible.
@@
Figure 2: Finite-time averages of the dimensionless variance $q(T)$ for $a=0.25$ with (a) $D=0.25$ for h=0.02, 0.03125, 0.05, 0.07, 0.1h=0.02,\,0.03125,\,0.05,\,0.07,\,0.1 and (b) $D=2$ for h=0.02, 0.03125, 0.1, 1, 2h=0.02,\,0.03125,\,0.1,\,1,\,2. In ergodic regimes, $q(T)\to 0$. At intermediate times, the decay follows a diffusive scaling $q(T)\sim T^{-1/2}$, while at longer times—when accessible—a crossover to a faster decay $q(T)\sim T^{-1}$ is observed.
@@ III.1 Finite Time Averages
Figure 3: The probability density function of the initial states for $a=0.25$ with (a) $D=0.25$, (b) $D=2$ for different energy densities $h$. The black dashed lines in both panels refer to a power law decay with an exponent $\gamma=2$ ($\sim 1/{\tau_{+}}^{2}$), to point out the results that has $\gamma\leq 2$ (nonergodic), and results that has $\gamma>2$ (ergodic). Insets: tail exponent $\gamma(h)$ obtained from the local slope of the log-log PDF. Red circles show the mean of $\gamma(\tau_{+})$ over the last two decades and error bars its standard deviation.
@@ III.2 PDF of Excursion Times
As in the previous section, we choose the local norm densities, $a_{\ell}=|b_{\ell}|^{2}$, as our local observables. The diagnostic developed herein proceeds by calculating the mean norm density $a$ for the entire system and then tracking each $a_{\ell}(t)$. We count the instances when $a_{\ell}(t)$ crosses the equilibrium value (or so-called the ensemble average of the observables $\langle a_{\ell}\rangle_{\Gamma}=a$) and measure the time intervals between these piercings, which are called excursion times [9, 31]. Specifically, we classify the excursion times $\tau$ based on whether $a_{\ell}>a$ or $a_{\ell}<a$, denoting them as $\tau_{+}$ and $\tau_{-}$, respectively. We then determine the probability distribution function (PDF) for $\tau_{+}$ and $\tau_{-}$. To compute $\mathrm{PDF}(\tau_{\pm})$, we bin the measured $\tau_{\pm}$ values into 200 bins and normalize by the total number of events and the bin width.
@@
where $PDF(\tau)$ denotes either $PDF(\tau_{+})$ or $PDF(\tau_{-})$, $\tau_{0}>0$ is a lower cutoff (a minimal excursion time), and the power-law form is assumed for the tail of $PDF(\tau)$. If the power-law decay exponent $\gamma\leq 2$, the mean of $\tau$ diverges, which signifies infinitely large excursion times. Thus, we expect $\gamma\leq 2$ for a nonergodic system (see [9, 31] for details).
@@
For each initial state (see Fig. 3), we extract the tail exponent from the local slope of the log–log representation of $\mathrm{PDF}(\tau_{+})$. The local exponent is defined as
$\gamma(\tau_{+})\equiv-\,\frac{d\log\mathrm{PDF}(\tau_{+})}{d\log\tau_{+}}.$ (20)
To suppress tail fluctuations, the $\log\mathrm{PDF}$ is first smoothed using a Hodrick–Prescott filter, yielding $\log\widetilde{\mathrm{PDF}}$. The derivative is then estimated by a centered finite difference along $\log\tau_{+}$,
$\gamma_{i}\approx-\frac{\log\widetilde{\mathrm{PDF}}_{i+1}-\log\widetilde{\mathrm{PDF}}_{i-1}}{\log\tau_{+,i+1}-\log\tau_{+,i-1}}.$ (21)
The analysis is restricted to the asymptotic tail region $\tau_{+}\in[10^{-2}\tau_{\max},\,\tau_{\max}]$. The reported value of $\gamma$ is the mean of $\gamma(\tau_{+})$ over this window, while the error bar represents its standard deviation, reflecting residual tail fluctuations rather than fitting uncertainty.
@@ III.3 Mechanism for few-site localization
3 sites. $S=\{\ell_{0}-1,\ell_{0},\ell_{0}+1\}$, set $A_{\ell_{0}-1}=A_{\ell_{0}+1}=x$, $A_{\ell_{0}}=y$, with $2x+y=A$. The reduced energy is
$H(x,y)=\frac{1}{2}x^{2}+\frac{1}{4}y^{2}-D\,xy\cos(2\Delta\phi).$
For the staggered compacton ($\Delta\phi=\pi/2$), a constrained extremization yields
$H_{3,\mathrm{stag}}(D)=\frac{1-2D^{2}}{4(3-4D)}\,A^{2},$ (24)
with the derivation given in Appendix B. For the in-phase configuration ($\Delta\phi=0$), the analogous three-site calculation gives $x=\frac{1+D}{4D+3}\,A$, $y=\frac{1+2D}{4D+3}\,A$, and
$H_{3,\mathrm{in}}(D)=\frac{1-2D^{2}}{4(4D+3)}\,A^{2}.$ (25)
In particular, for $D=2$ one finds $H_{3,\mathrm{in}}=-7A^{2}/44$, which is the global minimum at fixed $A$; cf. also the associated discussions of [29] and [35].
From Eqs. (22)–(25), we obtain the energy ordering at fixed $A$:
@@
Hence, for $D=0.25$ the constrained maximizer is the 1-site state. For large $H$ at fixed $A$, the dynamics select this constrained maximizer for $D<1$ [35], consistent with our $a_{\ell}(t)$ results (see Figs. 4,6). For $D=2$, the constrained maximizer is $H_{2,\mathrm{stag}}$ (with $H_{3,\mathrm{stag}}$ slightly lower), in agreement with Eq. (27), while the global minimum is attained by the three-site in-phase state, $H_{3,\mathrm{in}}(2)=-7A^{2}/44$ [cf. Eq. (25)].
For $D>1$, the constrained maximizer in the staggered family is the two-site staggered compacton [35], which dominates at large $H$ for fixed $A$. It is thus relevant to appreciate that the predominant dynamical observation in the case of $D=2$ involves such a 2-site staggered configuration, rather than the three-site states which are not generic in this case. Indeed, in our tests the latter appear only when the initial data excite two or more adjacent sites with nearly equal phases (see Eq.(15) as an example initial state used for such observation).
@@
Figure 5: High-energy localization on two and three sites for $D=2$. Local norm densities $a_{\ell}$ are color-coded in time for $a=0.25,h=1$. (a) Full lattice. (b) Zoom of (a) showing a 3-site, in-phase localization. (c) Zoom of (a) showing the dominant 2-site staggered localization. The initial state is prepared according to Eq. (15) with $a=0.25$, $x=19$, and $c=21$.
Figure 6: Nonergodic dynamics exhibits a one-site compacton for $D<1$ and a two-site staggered compacton for $D>1$. Local norm densities are shown as a function of time for $a=0.25$: (a) $D=0.75$, $h=0.57$ and (b) $D=1.25$, $h=1.25$. The initial states are prepared using Eq. (14) with the same random realization of $\eta_{\ell}$ and $\epsilon_{\ell}$, using $(c,x)=(4,12)$ in (a) and $(c,x)=(10,18.13)$ in (b). We plot only a subset of lattice sites (a zoom in $\ell$) to highlight the localization pattern.
@@ IV Discussion
For $D=0.25$, the system shows predominantly 1-site localization (see Fig. 4), whereas for $D=2$, we observed 2-site and 3-site stable localization (see Fig. 5). The 2-site (staggered) structures appear spontaneously and are generically favored, although, as is shown in the figure, an initialization of a 3-site in-phase excitation yields a structure that is quite long-lived, suggesting connections with the work of [29] on energy minimizers. These findings are consistent with previous work [35], which identified multi-site compactons in similar nonlinear systems. The behavior of the toy model for different values of $D$ correlates well with the theoretical predictions of localization patterns based on energy maximizers.
@@ Acknowledgments
This work was funded by the Deutsche Forschungsgemeinschaft (DFG) – Project-ID 414984028 – SFB 1404 FONDA and supported through the DFG Walter Benjamin Programme (Grant No. 506198590) (Y.K.). We acknowledge support from the Ministry of Science, Technological Development and Innovation of the Republic of Serbia, Grant No. 451-03-33/2026-03/200017, 451-03-34/2026-03/200124. This material is based upon work supported by the US National Science Foundation under Grant No. PHY-2110030, PHY-2408988 and DMS-2204702 (P.G.K.). This research was partly conducted while P.G.K. was visiting the Okinawa Institute of Science and Technology (OIST) through the Theoretical Sciences Visiting Program (TSVP) and the Sydney Mathematical Research Institute (SMRI) of the University of Sydney. This work was also supported by the Simons Foundation [SFI-MPS-SFM-00011048, P.G.K.]. The authors also acknowledge constructive discussions with Professor Jeremy Marzuola in connection to the work of [29] and its connections to the present manuscript.
@@
The numerical datasets used to generate the figures in this work are deposited on Zenodo with restricted access (DOI: 10.5281/zenodo.18682230). Access prior to public release is available from the corresponding author upon reasonable request.
@@ Appendix A Phase Diagram for Toy Model
Infinite temperature, $\beta\to 0$
We consider the infinite-temperature limit $\beta\to 0$ with
$\mu:=\beta\alpha\quad\text{fixed}\qquad(\alpha\to\infty).$ (28)
Starting from the partition function in Eq. (II), we use $I_{0}(0)=1$ and obtain, to leading order in $\beta$,
$Z\!\simeq\!(2\pi)^{N}\!\prod_{\ell=1}^{N}\int_{0}^{\infty}\!\!\!dA_{\ell}\exp\!\left[\!-\mu A_{\ell}\!-\!\frac{\beta}{4}A_{\ell}^{2}\right]\!\!=\!(2\pi)^{N}J(\beta,\mu)^{N},$ (29)